# A Thorough Assessment of the Non-IID Data Impact in Federated Learning

### v1

@@ Abstract
Federated learning (FL) allows collaborative machine learning (ML) model training among decentralized clients’ information, ensuring data privacy. The decentralized nature of FL deals with non-independent and identically distributed (non-IID) data. This open problem has notable consequences, such as decreased model performance and larger convergence times. Despite its importance, experimental studies systematically addressing all types of data heterogeneity (a.k.a. non-IIDness) remain scarce. This paper aims to fill this gap by assessing and quantifying the non-IID effect through a thorough empirical analysis. We use the Hellinger Distance (HD) to measure differences in distribution among clients. Our study benchmarks five state-of-the-art strategies for handling non-IID data, including label, feature, quantity, and spatiotemporal skewness, under realistic and controlled conditions. This is the first comprehensive analysis of the spatiotemporal skew effect in FL. Our findings highlight the significant impact of label and spatiotemporal skew non-IID types on FL model performance, with notable performance drops occurring at specific HD thresholds. Additionally, the FL performance is heavily affected mainly when the non-IIDness is extreme. Thus, we provide recommendations for FL research to tackle data heterogeneity effectively. Our work represents the most extensive examination of non-IIDness in FL, offering a robust foundation for future research.
Index Terms:
Federated Learning, Machine Learning, non-IID data, data heterogeneity quantification.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
In today’s digital age, the interaction of machine learning (ML) and healthcare or financial data holds immense promise for improving disease diagnosis [36] and combating financial crimes [35]. However, it raises critical questions about data privacy when dealing with sensitive information from hospitals or banks. In this context, trusted research environments emerge as a mechanism for balancing ML research and protecting individual privacy [13].
@@
Federated learning (FL) [30] has emerged as a transformative approach for training ML models across decentralized data sources, preserving data privacy and security. However, a significant challenge inherent in FL is the variation in data distributions across clients referred to as non-IID (Non-Independent and Identically Distributed) data. This so-called non-IIDness hinders model performance and convergence during training [44].
@@
Despite the growing body of research focusing on the challenges posed by non-IID data in FL, there remains a gap in experimental studies that systematically evaluate the effectiveness of FL algorithms under different types of data heterogeneity. Understanding the extent and nature of non-IID characteristics are crucial for developing robust, fair, and widely applicable FL algorithms. In this context, our study uses the Hellinger Distance (HD[12] to quantify the differences in client data distributions. This approach ensures that our conclusions are robust and generalizable across diverse scenarios.
@@
    We benchmark five of the most employed state-of-the-art aggregation and client selection of FL algorithms to tackle non-IID data distributions among clients under realistic, controlled, and quantifiable methods for synthetic partitioning and all non-IID types (label, feature, quantity, and spatiotemporal skewness). This is the first study thoroughly analyzing how the spatiotemporal skew affects the performance of FL models.
@@
    We motivate using HD to quantify the differences among data distributions, standardizing the guidelines for systematic studies of data heterogeneity in FL. This is the first work to demonstrate that the effect of the non-IIDness is not the same under all the levels of heterogeneity We use HD to quantify differences in distribution as it provides more granular information, and we leave as future work the exploration of other measures; see our section on design insights and opportunities.
@@
    We provide a reference to researchers about which methods are robust to which kind of non-IIDness on highly benchmarked datasets.
@@
    We give highlights and relevant recommendations for FL researchers based on quantifying non-IIDness.
@@
To the best of our knowledge, this is the most comprehensive and complete study of non-IIDness and its effects on FL models.
II Related Work
@@
Studies that analyze and benchmark the performance of methods to tackle the non-IIDness effect on the FL models under controlled and systematic scenarios are scarce. Nevertheless, in this section, we introduce those works that, to some extent, provide empirical analysis about the data heterogeneity repercussions.
@@
A study by Vahidian et al. [43] challenges conventional thinking regarding data heterogeneity in FL. They posit that dissimilar data among participants is not always detrimental and can be advantageous, and we found similar results. Their argument centers on two main points: firstly, that differences in labels (label skew) are not the sole determinant of data heterogeneity, and secondly, that a more effective measure of heterogeneity is the angle between the data subspaces of participating clients. Conversely, we encompass a broader spectrum of non-IIDness types and include images and tabular data.
@@
Wong et al. [45] conduct extensive experiments on a large network of IoT and edge devices to present FL real-world characteristics, including learning performance and operation (computation and communication) costs. Moreover, they mainly concentrate on heterogeneous scenarios, which is the most challenging issue of FL. Although they thoroughly analyze the impact of non-IIDness, they focus on image datasets and do not compare any aggregation method to solve the effects of highly heterogeneous scenarios. Contrarily, this work includes diverse datasets and benchmark state-of-the-art methodologies to handle non-IIDness in FL.
@@
In their study, Mora et al. [31] examine existing solutions in the literature to mitigate the challenges posed by non-IID data. On the one hand, they emphasize the underlying rationale behind these alternative strategies and discuss their potential limitations. Ultimately, they identify the most promising approaches based on empirical results and critical defining characteristics, such as any assumptions made by each strategy. In parallel, they focus only on label skew and only consider one dataset in their experiments. In our paper, we alternatively analyze broader datasets and data skewness types, identifying broader limitations and potential approaches to overcome them.
Wong et al. [45] present a comprehensive analysis of FL characteristics in real-world environments, focusing on large networks of IoT and edge devices. Their work offers valuable insights into learning performance and operational costs, particularly in heterogeneous scenarios. While they provide a detailed exploration of non-IID data issues using image datasets, they do not extensively evaluate different aggregation methods to address highly heterogeneous scenarios. In contrast, our work broadens this scope by incorporating diverse datasets and benchmarking state-of-the-art methods specifically designed to tackle non-IID challenges in FL.
One of the novelties of our study is the use of metrics to quantify non-IIDness among clients. The latter facilitates the systematic and regulated selection of non-IID. To the best of our knowledge, our work is the first to employ these metrics in the relevant bibliography.
III Background
III-A Federated learning basics
@@
FL [30] is particularly suited for scenarios with siloed data (clients, local nodes, parties) where multiple organizations or institutions hold their datasets. This decentralized approach enables collaborative model training without sharing the raw data, contributing to data privacy and ownership for each participating organization [10].
@@
Figure 1 presents an overview of the cross-silo FL training process, where participating clients train local models ($M_{i}$) on their datasets ($\mathcal{D}_{i}$), all based on a pre-distributed global model ($M$[38, 48]. Instead of sharing raw data, clients exchange model updates ($M_{i}$) without sensitive information, aggregating centrally to enhance the global model. By utilizing a central server for coordination, silos transmit their updates, aggregated to improve the global model. This iterative process allows collaboration without compromising data privacy, as each client receives the updated global model without sharing raw data, ensuring data privacy while facilitating collaboration.
Data Skewness Types.
A centralized dataset [^1] Notice that this definition of a centralized dataset includes tabular data, images, medical data, graph data, and any dataset expressible as a collection of arrays. $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),\dots,(\mathbf{x}_{n},y_{n})\}$, is a collection of $n$ tuples where $\mathbf{x}_{i}=[(x_{i})_{1},\dots,(x_{i})_{m}]$ is the feature representation of the $i$th element (sample) in the dataset, and $y_{i}\in\{1,\dots,\ell\}$ is the (true) label of the $i$th element.
@@
Defining the type of non-IIDness in FL is relevant since it can drastically influence the performance of the models. We follow the settings of previous work [28, 50]. For a supervised learning task on client $i$ (local node $i$), we assume that each data sample $(x,y)\in\mathcal{D}_{i}$, where $x$ is the input attributes or features, and $y$ is the label, following a local distribution $P_{i}(\mathbf{x},y)$. Let us define:
$P_{i}^{Y}(y)=\sum_{\begin{subarray}{c}(\mathbf{x},z)\in\mathcal{D}_{i}\\ z=y\end{subarray}}P_{i}(\mathbf{x},z)\hskip 8.53581pt$ $P_{i}^{X_{\ell}}(x)=\sum_{\begin{subarray}{c}(\mathbf{x},y)\in\mathcal{D}_{i}\\ x_{\ell}=x\end{subarray}}P_{i}(\mathbf{x},y)$
with $P_{i}^{Y}(y)$, the $i$th client labels’ distribution and $P_{i}^{X_{\ell}}(x)$ the distribution over the $\ell$th input feature of the $i$th client. Then, the classification for non-IID data (i.e., data skewness types) is as follows:
@@
      Spatiotemporal skewness: Refers to the inner correlation of data in the time (or space) domain. In other words, the distribution $P_{i}(\mathbf{x},y)$ is not stationary but depends on the time or space.
Quantifying the degree of non-IIDness.
Regarding the selection of valuable scenarios to demonstrate the effects of non-IIDness in FL, the current literature often relies on ad-hoc partitions [26, 21, 15]. Therefore, in this work, we use a metric that systematically evaluates the level of non-IID data to select scenarios for measuring the effect of non-IIDness. We opted for the Hellinger Distance (HD), a metric widely used to gauge the separation between two probability distributions calculated as in Equation 1 [12].
@@
Aggregation and client selection algorithms
@@
In an FL process, the server aggregates the weights obtained from each client and communicate them back to each participant. In this section, we explain the five state-of-the-art aggregation and client-selection algorithms assessed in our experiments.
@@
FedAvg: It is a fundamental algorithm in FL [30] designed to train ML models across a network of decentralized devices while preserving data privacy. In FedAvg, each client computes model updates and sends them to a central server using local data. The server averages these updates to calculate a global model update and then sends it back to the clients. The latter process iterates until it convergences (the model’s performance gets stable).
@@
FedProx: It is a framework designed to address heterogeneity in FL [24], offering a generalized and reparametrized version of FedAvg. This approach incorporates a regularization term ($\mu$) to minimize the difference between local and global weights. Finally, the framework aggregates local model updates from all devices to obtain an updated global model. Using this proximal term, FedProx aims to improve convergence and performance in heterogeneous federated learning environments.
@@
Random size-proportional selection (Rand): Rand is a baseline client selection strategy introduced to handle non-IID data that does not bias toward clients with higher local losses. Most current analysis frameworks consider a scheme that selects the training set of clients $S(t)$ by sampling $m$ clients randomly (with replacement) such that client $k$ gets selected with probability $p_{k}$, the fraction of data at that client [6].
@@
Power-Of-Choice (POC): The POC algorithm performs well under non-IID distribution. It is inspired by the power of $d$ choices load balancing strategy, which queueing systems commonly use [6][7]. The central server first samples a candidate set of $d$ clients, where $d$ is between $m$ (the number of clients to be selected) and $K$. These candidates are chosen based on their data fraction ($p_{k}$). The server then sends the current global model to these candidates, who compute and return their local losses. Finally, the server selects $m$ clients with the highest losses to participate in the next training round. This approach aims to balance the workload and prioritize clients with more informative updates, improving the efficiency of the FL process.
@@
Model Contrastive Learning (MOON): MOON is a simple and effective FL framework designed to tackle non-IIDness. It uses the similarity between model representations to correct the local training of individual parties (i.e., conducting contrastive learning at the model level) [23]. The network proposed in MOON has three components: a base encoder, a projection head, and an output layer. The base encoder extracts representation vectors from inputs. Le et al. introduce an additional projection head to map the representation to a space with a fixed dimension. Last, the output layer produces predicted values for each class. For ease of presentation, with model weight $w$, they use $F_{w}(\cdot)$ to denote the whole network and $R_{w}(\cdot)$ to denote the network before the output layer (i.e., $R_{w}(X_{\ell})$ is the mapped representation vector of input $X_{\ell}$).
IV Experimentation Setup
Datasets

### v2

@@ Abstract
Federated learning (FL) allows collaborative machine learning (ML) model training among decentralized clients’ information, ensuring data privacy. The decentralized nature of FL deals with non-independent and identically distributed (non-IID) data. This open problem has notable consequences, such as decreased model performance and longer convergence times. Despite its importance, experimental studies systematically addressing all types of data heterogeneity (a.k.a. non-IIDness) remain scarce. This paper aims to fill this gap by assessing and quantifying the non-IID effect through an empirical analysis. We use the Hellinger Distance (HD) to measure differences in distribution among clients. Our study benchmarks five state-of-the-art strategies for handling non-IID data, including label, feature, quantity, and spatiotemporal skews, under realistic and controlled conditions. This is the first comprehensive analysis of the spatiotemporal skew effect in FL. Our findings highlight the significant impact of label and spatiotemporal skew non-IID types on FL model performance, with notable performance drops occurring at specific HD thresholds. The FL performance is also heavily affected, mainly when the non-IIDness is extreme. Thus, we provide recommendations for FL research to tackle data heterogeneity effectively. Our work represents the most extensive examination of non-IIDness in FL, offering a robust foundation for future research.
keywords:
[…]
{highlights}
Label skew and spatiotemporal skew have the most significant impact on the model’s performance.
The drop in the model’s performance for label skew appears in a double threshold. A notable performance decline is immediately evident when the Hellinger Distance exceeds 0.5 and 0.75.
Feature skew does not alter the model’s performance nor the convergence point.
The quantity skew in the client’s data does not affect the model’s performance.
The higher the non-IIDness level in time and space (spatiotemporal skew), the worse the model’s performance.
1 Introduction
@@
In today’s digital age, the interaction of machine learning (ML) and healthcare or financial data holds immense promise for improving disease diagnosis [57] and combating financial crimes [55]. These advancements have traditionally relied on centralized learning (CL), such as Machine Learning as a Service (MLaaS) platforms—including AWS, Azure, and Google Cloud [2]—where raw data is aggregated on a central server for model training.. However, it raises critical questions about data privacy when dealing with sensitive information from hospitals or banks. In this context, trusted research environments emerge as a mechanism for balancing ML research and protecting individual privacy [23, 80].
@@
Federated learning (FL) [47] has emerged as a transformative approach for training ML models across decentralized data sources, preserving data privacy and security. This paradigm is particularly beneficial in cross-silo settings, where entities such as hospitals, banks, and other organizations collaborate without sharing sensitive data. However, a significant challenge inherent in FL is the variation in data distributions across clients, referred to as non-IID (Non-Independent and Identically Distributed) data. This non-IID data (i.e., non-IIDness, data heterogeneity) hinders model performance and convergence during training [66, 48]. Such non-IID data is classified into four categories: label, feature, quantity, and spatiotemporal skew [83].
Spatiotemporal skew presents unique challenges that are particularly critical yet underexplored in FL research. This skew occurs when data distributions vary across both geographical locations (spatial) and periods (temporal) [15]. Such skew fundamentally differs from the label, feature, or quantity skew by introducing dynamic variations that standard FL aggregation algorithms often fail to address [4]. Understanding spatiotemporal skew is crucial because it directly impacts the model’s ability to generalize across diverse real-world environments while maintaining temporal relevance, making it a key frontier for robust FL systems.
Furthermore, diagnosing and quantifying the level of non-IID data in FL is a significant challenge, as emphasized by Pei et al. [56] and Li et al. [41], who identify critical research directions in this domain. Numerous studies have introduced metrics to quantify the level of non-IID data in FL [33, 19, 52, 67], with the Hellinger Distance (HD[22] emerging as one of the most reliable options. HD = 0.0 corresponds to fully IID data, while higher values (e.g., 0.25, 0.5, 0.75, 0.9) represent increasing degrees of non-IID data, with HD = 0.9 approaching the most non-IID scenario considered in our study. HD offers a fine-grained measurement of distribution differences, achieving values close to 1 under extreme non-IID conditions, unlike the Jensen-Shannon Distance (JSD), which tends to plateau at lower levels. Furthermore, HD is versatile and applicable across various types of skews.
@@
Recent advances in FL research have significantly advanced our understanding of non-IID data challenges, with notable progress in addressing isolated aspects of heterogeneity such as label skew [64, 69, 49]. Based on this foundation, there is now a timely opportunity to unify these insights through comprehensive empirical benchmarks that span the full spectrum of non-IID skews. While pioneering theoretical frameworks [44, 56] have established critical mitigation principles, translating these into practice requires systematic quantification of how diverse non-IID data types—from feature skew to spatiotemporal drift—affect real-world FL performance metrics. Closing this knowledge gap through rigorous experimental analysis will empower the community to develop FL systems that are theoretically sound and empirically robust across application domains.
Our study addresses this gap by using the HD to quantify differences in client data distributions, enabling a rigorous empirical analysis of non-IID effects across multiple dimensions, including label, feature, quantity, and spatiotemporal skews. Throughout the assessment of spatiotemporal skew, we capture the impact of dynamic data shifts over time and space, which are particularly relevant in real-world applications such as banking credit risk [81] and personalized healthcare [25]. This approach ensures that our conclusions are robust and generalizable across diverse scenarios.
@@
    We benchmark five of the most employed state-of-the-art aggregation and client selection of FL algorithms to tackle non-IID data distributions among clients under realistic, controlled, and quantifiable methods for synthetic data partitioning and all non-IID types (label, feature, quantity, and spatiotemporal skews). Previous empirical works have focused on label skew [64, 69, 49] or in combinations of label, feature, and quantity skew [38] (see Section2.1 for more details). Thus, this is the first study empirically analyzing how the spatiotemporal skew affects the performance of FL models.
@@
    We motivate using HD to quantify the differences among data distributions, standardizing the guidelines for systematic studies of non-IID data in FL. This is the first work to demonstrate that the effect of the non-IID data is not the same under all levels of heterogeneity. We use HD to quantify differences in distribution as it provides more granular information, and we leave as future work the exploration of other measures; see our section on design insights and opportunities.
@@
    We provide a reference to researchers about which methods are robust to which kind of non-IID data on highly benchmarked datasets.
@@
    We give highlights and relevant recommendations for FL researchers based on quantifying the level of non-IID data.
@@
To the best of our knowledge, this is the most comprehensive and complete empirical study of non-IID data and its effects on FL models.
Positioning within Industrial Information Integration.
This study contributes to the Journal of Industrial Information Integration’s focus on industrial non-IID data and privacy-preserving analytics by thoroughly evaluating non-IID data effects in FL. FL is increasingly adopted in industrial domains such as healthcare, intrusion detection in IoT, and digital twins for industrial IoT, where data is distributed across multiple silos and devices with inherent heterogeneity. Prior works in this journal have addressed related challenges, including emotion recognition based on electroencephalography (EEG) as a crucial research area in the Internet of Medical Things (IoMT) [32], federated ensemble model for intrusion detection in distributed IoT networks for enhancing cybersecurity [9], and adaptive optimization for FL enabled digital twins in industrial IoT [73].
Our work advances these efforts by systematically quantifying the impact of diverse non-IID data types on FL model performance using the HD metric. Furthermore, we benchmark state-of-the-art aggregation and client selection algorithms, offering practical guidance for deploying FL in industrial scenarios characterized by complex data distributions. This positioning situates our research within the ongoing scholarly conversation on industrial information integration and FL, underscoring its significance for robust, privacy-aware industrial analytics.
Ethical Considerations in FL:.
FL inherently aligns with ethical principles related to data minimization and user privacy, as it allows individual clients to retain their raw data locally. However, despite these benefits, FL is not immune to ethical concerns. Potential privacy risks remain due to model inversion or gradient leakage attacks [27], and there is a need for transparency and informed consent when deploying FL in real-world applications [54]. Future work must integrate privacy-preserving mechanisms (e.g., differential privacy [68], secure aggregation [20]) and conduct rigorous audits to ensure ethical compliance in decentralized learning scenarios [77].
2 Related Work
In this section, we present recent studies that empirically evaluate the behavior of non-IID data in FL. Additionally, we compare our work with relevant surveys on non-IID data in FL.
2.1 Empirical Studies
@@
Studies that analyze and benchmark the performance of methods to tackle the non-IID data effect on the FL models under controlled and systematic scenarios are scarce. Nevertheless, in this section, we introduce those works that, to some extent, provide empirical analysis about the repercussions of non-IID data.
@@
A study by Vahidian et al. [64] challenges conventional thinking regarding non-IID data in FL. They posit that dissimilar data among participants is not always detrimental and can be advantageous, and we found similar results. Their argument centers on two main points: firstly, that differences in labels (label skew) are not the sole determinant of non-IID data, and secondly, that a more effective measure of heterogeneity is the angle between the data subspaces of participating clients. Complementary, we encompass a broader spectrum of non-IID data types and include images and tabular data.
@@
Wong et al. [69] conduct extensive experiments on a large network of IoT and edge devices to present FL real-world characteristics, including learning performance and operation (computation and communication) costs. Moreover, they mainly concentrate on heterogeneous scenarios, the most challenging issue of FL. While they thoroughly analyze the impact of non-IID data, the focus is primarily on image datasets, and they do not explore comparisons of aggregation algorithms to address highly heterogeneous scenarios. In contrast, our work expands on this by incorporating diverse datasets and benchmarking state-of-the-art methodologies to tackle non-IID data in FL effectively, offering a broader and more practical perspective.
@@
In their study, Mora et al. [49] examine existing solutions in the literature to mitigate the challenges posed by non-IID data. On the one hand, they emphasize the underlying rationale behind these alternative strategies and discuss their potential limitations. On the other hand, they identify the most promising approaches based on empirical results and critical defining characteristics, such as any assumptions made by each strategy. In addition, they focused on label skew and considered one dataset in their experiments. In our paper, we alternatively analyze broader datasets and data skewness types, identifying limitations and potential approaches to overcome them.
Li et al. [38] conduct a comprehensive experimental evaluation of FL aggregation algorithms under non-IID data settings. They systematically analyze the strengths and limitations of state-of-the-art FL aggregation algorithms while introducing diverse data partitioning methods to simulate various non-IID scenarios. Their work highlights the challenges posed by non-IID data, such as accuracy degradation and training instability, and provides empirical insights into the performance of each algorithm across different settings. In our paper, we build upon their work by analyzing spatiotemporal skew and introducing metrics to quantify its effects, offering a broader perspective on non-IID data and its impact on model performance.
2.2 Surveys
Some surveys that explore the effects of non-IID data in the model performance have been proposed, and they are depicted in Table 1. Earlier works, such as those from 2024, focus on label skew, providing valuable insights into this dimension. Resources from 2022 partially mention spatiotemporal skew, contributing to a deeper understanding of these aspects. The 2021 study offers an essential foundation by addressing label skew, paving the way for more comprehensive analyses in subsequent years.
Resource Publication Year Label Skew Feature Skew Quantity Skew Spatiotemporal Skew Non-IID data Quantification Empirical Highlights
[…]
Compared to the previous surveys summarized in Table 1, our work builds upon and extends their contributions by offering a more comprehensive approach. We include empirical evaluations of the effects of non-IID data, provide quantification of the non-IID data level, and conduct extensive experiments to assess the impact of spatiotemporal skew thoroughly. These additions enhance our study’s depth and practical relevance, complementing earlier works.
3 Background
This section provides an overview of FL, its fundamental training process, and the challenges posed by non-IID data. We categorize different types of data skew that impact FL performance and introduce methods for quantifying the level of non-IID data. Additionally, we review state-of-the-art aggregation and client selection strategies designed to address these challenges, such as FedAvg [47], FedProx [41], Random size-proportional selection (Rand) [13], Power-Of-Choice (POC) [13], and Model Contrastive Learning (MOON) [40].
3.1 Basics of FL
@@
FL [47] suits siloed data (a.k.a. clients, local nodes, parties, participants) where multiple organizations or institutions hold their datasets. This decentralized approach enables collaborative model training without sharing the raw data, contributing to data privacy and ownership for each participating organization [18].
@@
Figure 1 presents an overview of the cross-silo FL training process, where participating clients train local models ($M_{i}$) on their datasets ($\mathcal{D}_{i}$), all based on a pre-distributed global model ($M$[59, 79]. Instead of sharing raw data, clients exchange model updates ($M_{i}$) without sensitive information, aggregating centrally to enhance the global model. By utilizing a central server for coordination, clients transmit their updates, aggregated to improve the global model. This iterative process allows collaboration without compromising data privacy, as each client receives the updated global model without sharing raw data, ensuring data privacy while facilitating collaboration.
3.2 Data skew types.
A centralized dataset [^1] Notice that this definition of a centralized dataset includes tabular data, images, medical data, and graph data, and any dataset expressible as a collection of arrays. $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),\dots,(\mathbf{x}_{n},y_{n})\}$, is a collection of $n$ tuples where $\mathbf{x}_{i}=[(x_{i})_{1},\dots,(x_{i})_{m}]$ is the feature representation of the $i$th element (sample) in the dataset, and $y_{i}\in\{1,\dots,\ell\}$ is the (true) label of the $i$th element.
@@
Defining the type of non-IID data in FL is relevant since it can drastically influence the performance of the models. We follow the settings of previous work [46, 83]. For a supervised learning task on client $i$ (local node $i$), we assume that each data sample $(x,y)\in\mathcal{D}_{i}$, where $x$ is the input attributes or features, and $y$ is the label, following a local distribution $P_{i}(\mathbf{x},y)$. Let us define:
$P_{i}^{Y}(y)=\sum_{\begin{subarray}{c}(\mathbf{x},z)\in\mathcal{D}_{i}\!\!\\ z=y\end{subarray}}P_{i}(\mathbf{x},z)\quad\text{and}\quad P_{i}^{X_{\ell}}(x)=\sum_{\begin{subarray}{c}(\mathbf{x},y)\in\mathcal{D}_{i}\!\!\\ x_{\ell}=x\end{subarray}}P_{i}(\mathbf{x},y)$ (1)
with $P_{i}^{Y}(y)$, the $i$th client labels’ distribution and $P_{i}^{X_{\ell}}(x)$ the distribution over the $\ell$th input feature of the $i$th client. Then, the classification for non-IID data (i.e., data skew types) is as follows:
@@
      Spatiotemporal skew: Also known as spatial-temporal skew under federated continual learning (FCL) [75, 74, 76]. It refers to the inner correlation of data in the time (or space) domain. In other words, the distribution $P_{i}(\mathbf{x},y)$ is not stationary but depends on time or space.
3.3 Quantifying the Degree of non-IID data.
Regarding selecting valuable scenarios to demonstrate the effects of non-IID data in FL, the current literature often relies on ad-hoc partitions [43, 39, 30]. Therefore, in this work, we use a metric that systematically evaluates the level of non-IID data to select scenarios for measuring the effect of non-IID data. We opted for the Hellinger Distance (HD), a metric widely used to gauge the separation between two probability distributions calculated as in Equation 2 [22].
@@
HD provides a fine-grained and sensitive measurement of distributional differences, reaching values close to 1 under extreme non-IID conditions—unlike JSD, which often saturates and fails to reflect high levels of skew. Additionally, HD is highly adaptable across different types of non-IID data. In contrast to Earth Mover’s Distance (EMD) [33], whose values depend heavily on the choice and scale of the ground distance, HD offers normalized and consistent comparisons across tasks and datasets, making it particularly well-suited for FL scenarios.
3.4 Aggregation and Client Selection Algorithms
@@
In an FL process, the server aggregates the weights obtained from each client and communicate them back to each participant.
In this section, we explain the five state-of-the-art aggregation and client-selection algorithms assessed in our experiments.
@@
FedAvg: It is a fundamental algorithm in FL [47] designed to train ML models across a network of decentralized devices while preserving data privacy. In FedAvg, each client computes model updates and sends them to a central server using local data. The server averages these updates to calculate a global model update and then sends it back to the clients. This process iterates until it converges (the model’s performance gets stable). FedAvg suffers from three key limitations: (1) degraded performance under non-IID data distributions across clients, (2) inefficient communication rounds caused by straggling devices or disproportionate local dataset sizes, and (3) a uniform aggregation approach that fails to account for variability in client data quality or device reliability, potentially biasing the global model.
@@
FedProx: It is a framework designed to address non-IID data in FL [41], offering a generalized and reparametrized version of FedAvg. This approach incorporates a regularization term ($\mu$) to minimize the difference between local and global weights. Finally, the framework aggregates local model updates from all devices to obtain an updated global model. Using this proximal term, FedProx aims to improve convergence and performance in heterogeneous federated learning environments. FedProx ensures convergence even with non-IID data while requiring only minor adjustments to implementation. However, it has notable drawbacks: (1) its effectiveness depends significantly on careful hyperparameter selection, and incorrect settings may slow convergence or raise communication overhead; (2) in real-world applications, its advantages weaken under extreme levels of non-IID data, particularly when client datasets contain completely distinct classes. [38].
@@
Rand: Rand is a baseline client selection strategy introduced to handle non-IID data that is not biased toward clients with higher local losses. Most current analysis frameworks consider a scheme that selects the training set of clients $S(t)$ by sampling $m$ clients randomly (with replacement) such that client $k$ gets selected with probability $p_{k}$, the fraction of data at that client [12]. Rand provides unbiased client selection but suffers from key limitations: (1) it fails to prioritize clients with informative updates, hindering convergence speed; and (2) in highly non-IID settings, it may underrepresent rare data distributions, reducing model generalization.
@@
POC: The POC algorithm performs well under a non-IID distribution. It is inspired by the power of $d$ choices load balancing strategy, which queueing systems commonly use [13, 12]. The central server first samples a candidate set of $d$ clients, where $d$ is between $m$ (the number of clients to be selected) and $K$. These candidates are chosen based on their data fraction ($p_{k}$). The server then sends the current global model to these candidates, who compute and return their local losses. Finally, the server selects $m$ clients with the highest losses to participate in the next training round. This approach aims to balance the workload and prioritize clients with more informative updates, improving the efficiency of the FL process. While this approach enhances training efficiency and manages non-IID data effectively, it also has limitations: (1) Selection bias can marginalize underrepresented clients, thereby weakening the model’s generalization ability. (2) The method introduces higher complexity and greater communication costs in the client selection process.
@@
MOON: It is a simple and effective FL framework designed to tackle non-IID data. It uses the similarity between model representations to correct the local training of individual parties (i.e., conducting contrastive learning at the model level). The network proposed in MOON has three components: a base encoder, a projection head, and an output layer. The base encoder extracts representation vectors from inputs. Le et al. [40] introduce an additional projection head to map the representation to a space with a fixed dimension. Last, the output layer produces predicted values for each class. For ease of presentation, with model weight $w$, they use $F_{w}(\cdot)$ to denote the whole network and $R_{w}(\cdot)$ to denote the network before the output layer (i.e., $R_{w}(X_{\ell})$ is the mapped representation vector of input $X_{\ell}$). Like the previous methods, this approach is effective but has drawbacks: (1) higher computational costs due to generating and comparing augmented data views, and (2) restricted use for non-visual data (e.g., text or time-series), where creating meaningful augmentations is difficult because of text context-dependence and time-series structural limitations. [11]
3.5 Models used in FL
In FL, the choice of model architecture plays a critical role in determining both performance and communication efficiency across distributed clients. Depending on the nature of the data and the target task, different types of models may be employed to balance expressiveness, computational cost, and generalizability [70]. Below, we outline several common model types used in FL, highlighting their core characteristics and suitability for decentralized training environments.
  • Deep Neural Networks (DNNs): DNNs are feedforward networks with multiple hidden layers, capable of learning complex patterns through hierarchical feature extraction [3]. They serve as foundational models in FL due to their flexibility.
  • Convolutional Neural Networks (CNNs): CNNs specialize in processing grid-like data (e.g., images) using convolutional layers for local feature detection, pooling for dimensionality reduction, and fully connected layers for classification [28]. Their parameter-sharing property makes them efficient for FL tasks.
  • Transfer Learning Models: Pre-trained architectures like ResNet9 [29], EfficientNetB0 [63], and MobileNetV2 [60], leverage transfer learning by adapting learned features from large datasets (e.g., ImageNet) to new tasks with limited data [65]. In FL, such models reduce communication overhead and improve convergence by starting from robust initial weights.
4 Experimentation Setup
Datasets.